# Semiparametric Discovery and Estimation of Interaction in Mixed Exposures using Stochastic Interventions

### v2

@@ 3.2 TMLE
We chose to estimate our parameter using TMLE (rather than estimating equations, inverse propensity of treatment weights (IPTW) and augmented IPTW ) chiefly due to its favorable finite sample properties due to being a plug-in estimator. Briefly, to construct the TMLE estimator, we: 1. Use data-adaptive regression for the initial estimation of $g_{0}(\mathbf{a},W)$ and $\overline{Q}_{0}(\mathbf{a},W)$, 2. Evaluate $H(\mathbf{a}_{i},w_{i})$ for each observation, 3. Using the derived estimates, we implement a logistic regression to refine predictions under the shift $\text{logit}\overline{Q}_{\epsilon,n}(\mathbf{a}-\delta,w)=\text{logit}\overline{Q}_{n}(\mathbf{a}-\delta,w)+\epsilon H_{n}(\mathbf{a}-\delta,w)$ then 4. From this model, once $\epsilon$ is determined, we update the predicted counterfactuals, using these updated counterfactuals, we then compute the plugin. This is a standard TMLE procedure for stochastic interventions; we simply shift multiple exposures simultaneously in the initial estimation and update. We estimate a standard error by taking the empirical variance of the estimated EIF and use this construct Wald-type confidence intervals and p-values. We use TMLE to ascertain unbiased estimates for each part of our interaction target parameter. That is, in $E_{\boldsymbol{\delta}}[Y]$ we update our counterfactual outcome under a joint shift by the joint density ratio as $H(\boldsymbol{a},w)$. Similarly, $E_{\delta_{1}}[Y]$ and $E_{\delta_{2}}[Y]$ are simply updated based on $H(a,w)$ for univariate shifts. In general, we use TMLE to target each of the three parts of our parameter. Given that our interaction is a continuous function (linear combination) of these asymptotically efficient estimates, we use the delta method for variance derivation.
3.3 Joint Density Estimation
In the context of extending the clever covariate to account for joint shifts, a challenge emerges when confronting the estimation of the joint density, especially in scenarios with high-dimensional $W$. Direct estimation of joint conditional densities using conventional techniques scales poorly to large data Chen et al. [2018]. To avoid the complications inherent in the direct estimate of joint density, we take an alternative approach. Instead of seeking an explicit joint density estimator, the methodology takes advantage of the combination of conditional and marginal densities. In the bivariate scenario involving exposures $A1$ and $A2$, the joint density can be calculated more conveniently as $g_{0}(\mathbf{A}|W)=g_{0}(A1|W)\times g_{0}(A2|A1,W)$. Within this framework, two ensemble-based machine learning density estimators are deployed. The first is used to approximate the density of $A1$ conditioned on $W$, while the second estimates the conditional density of $A2$ given both $A1$ and $W$. The product of these constructed densities provides an effective estimate of the joint density, which avoids some of the inherent challenges of direct joint density estimation.
@@ 4.1 Identifying Interactions in the Mixture Using Training Data
The discovery of variables and interactions of mixture variables that are important for the prediction of the outcome is done using an ensemble method: the discrete Super Learner (SL) van der Laan Mark et al. [2007]. This algorithm employs a cross-validation mechanism, optimizing over a library of candidate algorithms to select the estimator with the best empirical fit. We use a discrete SL to select the best fitting estimator from a library of different sieve-based estimators that use basis splines and their tensor products, allowing the model to capture intricate nonlinear relationships as linear combinations of lower-dimensional basis functions while remaining interpretable. These include packages like earth Milborrow. Derived from mda:mars by T. Hastie and R. Tibshirani. [2011] , polySpline Ripley and Venables [2021], and hal9001 Coyle et al. [2022]. In our current implementation, our candidate algorithms used in the library are all based on earth. Because earth has many hyperparameters such as the number of interactions, how knot points are searched, etc. we include earth models with various parameters and select the best one based on cross-validation. This still falls within the framework of Super Learning. We employ this estimator for $E[Y|\boldsymbol{A},W]$ used in the data-adaptive discovery procedure using the training data. For example, for $k=1$ in a 10 fold CV, this includes all folds not equal to 1. In the context of high-dimensional mixture analysis, traditional ANOVA techniques are likely inadequate due to their limitations in handling nonlinear relationships. To address this, we introduce a non-parametric extension, Non-parametric Analysis of Variance (NP-ANOVA), following the selection of the best-fitting estimator through the Super Learner algorithm. The NP-ANOVA approach is designed to decompose the variance of the response variable by attributing it to contributions from individual basis functions utilized in the chosen model. This method is versatile and is applicable across various model structures such as MARS, HAL, among others. The central idea is to assess the relative importance of each basis function in explaining the variance of the response variable. The procedure can be summarized as follows. 1. From the model selected by the Super Learner, we extract the basis functions employed by the model, 2. Utilize these basis functions to build a GLM, allowing for a structured analysis of the contributions of bases, 3. Perform a Type 3 ANOVA on the GLM. This step involves systematically evaluating the change in explained variance when each basis function is omitted, thus assessing their individual contributions, 4. For each basis function, calculate an F-statistic, quantifying its impact on the explanatory power of the model, 5. Group and sum the F-statistics according to the variable sets corresponding to each basis, such as individual variables (e.g., $A_{1}$) or interactions (e.g., $A_{3}-A_{4}$), 6. Employ quantile-based thresholding on aggregated F-statistics to discern the most influential variable sets. For example, selecting the top 25% of variable sets or including all sets by setting the threshold at 0.
@@ 4.2 Estimation of Shift Interventions on Discovered Variable Sets
Treating these variable sets as fixed, we then estimate the resulting data-adaptively defined parameters using TMLE for marginal and joint stochastic shift interventions. Using the training data, the nuisance functions $g_{n}$ and $Q_{n}$ are trained. The TMLE update step is performed using these training estimates to estimate $\epsilon$. With the validation data, we obtain estimates for $g_{n}$ and $Q_{n}$ and apply $\epsilon$ to debias our initial plug-in for our estimated stochastic shift target parameter using the validation data. In our running example, the validation data is data that correspond to $k=1$. This procedure describes CV-TMLE Hubbard et al. [2018, 2016] which is the infrastructure we use to estimate our target parameter. Readers interested in the details of this procedure are directed to these resources.
@@
With the estimates for our parameter complete for the validation for $k=1$, we would then move on to $k=2$ and repeat the sections 2-4. Upon completion of the cross-estimation procedure, a pooled TMLE update provides a summary measure of the target parameter across k-folds. Specifically, we stack the predictions for each nuisance parameter from the estimation samples and run a pooled TMLE update on the cumulative initial estimates. The resulting average is then used for parameter estimation. The pooled estimate across the folds can be succinctly represented as $\theta=\frac{1}{K}\sum_{k}\Psi_{F_{T_{n,k}}}\left(V_{n,k}\right)(P_{0})$. Where $\Psi_{F_{T_{n,k}}}$ identifies optimal exposure sets and subsequently formulates a plug-in estimator using training data $T_{n,k}$. $V_{n,k}$ refers to our data’s estimation sample for deriving our estimates.
It is possible that the same interactions are not found in every fold. Here, our approach deviates from the standard data-adaptive target parameter method. In defining parameters normally, one would define, for example, the interaction with maximum impact defined by some measure (like our F-statistic or another measure) or perhaps create ranks. However, this may result in many different interactions being included in the top rank, which makes interpretation difficult. In this setting, the analyst is then required to interpret only the fold-specific results, which are statistically valid but have low power, and perhaps report the proportion of folds the interaction was found. In the proposed framework, we gain power by pooling in folds with consistent estimates, but the pooled result is difficult to interpret as well because these estimates do not incorporate information from folds with null results. As such, in refining our approach to pooled TMLE estimation, our goal is to integrate both the non-null estimates from the estimation samples and the null values. To address this, we propose a weighted average approach that combines our original pooled estimate with the implied null value, which we define as 0.
Specifically, let $n_{0}$ denote the number of folds where the null is implied, and let $n_{1}$ represent the number of folds with non-null estimates. The variance of the "null" estimate, $var(\Psi_{\text{null}})$, is estimated using the variance of the original pooled estimates, $var(\hat{\Psi}_{\text{pooled}})$, as follows $var(\Psi_{\text{null}})=\left(var(\hat{\Psi}_{\text{pooled}})\right)\cdot\frac{n_{1}}{n_{0}}$. We then calculate the weighted average of the original pooled estimate and the null value $\hat{\Psi}_{\text{pooled}}^{*}=\frac{w_{1}\cdot\hat{\Psi}_{\text{pooled}}+w_{0}\cdot\Psi_{\text{null}}}{w_{0}+w_{1}}$, where $w_{1}=\frac{1}{var(\hat{\Psi}_{\text{pooled}})}$ and $w_{0}=\frac{1}{var(\Psi_{\text{null}})}$. The variance of this new pooled estimate is given by $var(\hat{\Psi}_{\text{pooled}}^{*})=\frac{1}{w_{1}+w_{0}}$. This approach ensures that our pooled TMLE estimate reflects both the explicit estimates from folds where the variable set was selected and the implicit null values from the remaining folds. Of course, in settings where there is strong "signal" and interactions are consistently found across the folds this inverse variance approach is not needed and only the standard pooled TMLE is used. Fold-specific TMLE updates are also provided, allowing researchers to gauge consistency by comparing estimates across individual folds. Each fold-specific update applies a similar logic as the pooled update but remains confined to the respective fold, utilizing its specific $Q_{k}$ and $h_{k}$ parameters. Consequently, the EIF for each fold is calculated using only validation data within the fold, providing distinct variance estimates.
@@ 4.3 Pooled Estimates under Data-Adaptive Delta
Stochastic interventions, particularly joint stochastic shifts that involve products of conditional densities, may be vulnerable to positivity violations. Such violations can introduce bias and inflate variance in exposure effect estimation after shift, even with sophisticated estimators like TMLE. Addressing this necessitates a restrained $\mathbf{\delta}$, diminishing positivity violations, and thereby reducing bias and variance. We propose treating $\mathbf{\delta}$ as data-adaptive. It is determined in the training sample, adhering to positivity criteria, and subsequently applied to the validation sample. Define $H(a_{\delta},w)_{i}$ as the density ratio for observation $i$ after a shift $\delta$. To maintain ratios below a threshold $\lambda$, we iteratively diminish $\delta$ by $\epsilon$ until: $\forall i,\;H(a_{\delta},w)=\frac{g_{n_{-k}}(a_{n_{-k}}-\delta\mid w)}{g_{n_{-k}}(a_{n_{-k}}\mid w)}\leq\lambda$. Here, $\lambda$ is preset (default is 50 in our implementation). If any predicted conditional probabilities exceed this threshold under observed exposure, $\delta$ decreases by $\epsilon$ (defaulted to 10% in our package). This adjustment employs training data with estimators trained therein. Should $\delta$ remain constant across folds, the pooling involves averaging the estimates of the target parameters. For a data-adaptive $\delta$, where $\delta$ may be different in different folds, we simply take the mean estimate for a pooled $\delta$ across the folds that is paired with the pooled TMLE estimate.
@@ 5 Simulations
In this section, we demonstrate using simulations that our approach identifies the correct exposures and interactions used to generate the outcome and correctly estimates the counterfactual mean difference in outcome under joint shift intervention, marginal shifts, and subsequently the comparison, our interaction parameter. We construct a data-generating process (DGP) where $Y$ is generated from a linear combination of different effects including marginal impacts and interactions. This DGP was constructed to represent the complexity of a mixed exposure in which mixture variables are correlated but only some have an impact on the outcome.
This DGP has the following characteristics, $O=W,A,Y$. Let $W=(W_{1},W_{2},W_{3})$ denote a random vector of covariates, where $W_{1}$ and $W_{2}$ are bivariate normal with mean vector $\boldsymbol{\mu}=(6,7)$ and covariance matrix $\boldsymbol{\Sigma}_{W}=\begin{pmatrix}1&0.4\\ 0.4&1\end{pmatrix}$ and $W_{3}$ is a binary variable with Bernoulli distribution with probability $0.5$. The mean values for the components of the distribution of mixture $A$ are calculated as $\mu_{1}==e^{W_{1}/2}$, $\mu_{2}=W_{2}/2$, and $\mu_{3}=5$. That is, exposure 3 is not associated with the covariates. A exposure vector $A=(A_{1},A_{2},A_{3})$ is generated from a trivariate normal distribution with a mean vector $\boldsymbol{\mu}=(\mu_{1},\mu_{2},\mu_{3})$ and a covariance matrix: $\boldsymbol{\Sigma}_{A}=\begin{pmatrix}1&0.5&0.8\\ 0.5&1&0.7\\ 0.8&0.7&1\end{pmatrix}$. A fourth mixture variable was created via $A_{4}\sim\mathcal{N}(4,2)$. The outcome $Y$ is then generated by: $Y=1.3A_{4}+0.4A_{4}A_{1}+0.1W_{1}+0.3W_{2}+\epsilon+(A_{3}^{2}\mathbf{1}_{\{W_{3}=1\}}+A_{3}\mathbf{1}_{\{W_{3}=0\}})$. Where the third component of $A$ depends on $W_{3}$, $A_{3}$ is squared if $W_{3}$ is equal to 1, and is left unchanged if it is equal to 0. In generating the outcome, this variable represents effect modification or the differential impact of $A_{3}$ depending on $W_{3}$. Overall, built into this outcome-generating process there is a marginal impact of $A_{4}$, a multiplicative interaction between $A_{1}$ and $A_{4}$ and effect modification of $A_{3}$ based on $W_{3}$. There is no effect of $A_{2}$ on the outcome. This DGP was generated to reflect, although more simply, a real-world exposure situation that includes both interaction, effect modification, and no effects.
To calculate the ground-truth counterfactual mean under a shift intervention, we simply generate a very large data set from this DGP and apply various shifts to the exposures. We applied a shift of 1 unit increase to each exposure. In the first simulation, we deterministically set the variable sets to evaluate estimates of the counterfactual outcomes under these shifts. This was done with the var_sets parameter in SuperNOVA which, if is not null, bypasses the data-adaptive determination of the variable sets in the basis functions. In the second simulation we include the data-adaptive discovery of the variable sets to provide estimates for evaluating the variable relationship identification through the basis function procedure. In our case, given this DGP, we are interested in assessing the counterfactual mean differences under shifts for $A_{4}$ individually, $A_{1}$ individually, $A_{4}$ and $A_{1}$ jointly and comparing this joint shift with the individual shifts.

### v3

@@ 3.2 TMLE
We chose to estimate our parameter using TMLE (rather than estimating equations, inverse propensity of treatment weights (IPTW) and augmented IPTW ) chiefly due to its favorable finite sample properties due to being a plug-in estimator Li et al. [2022]. Briefly, to construct the TMLE estimator, we: 1. Use data-adaptive regression for the initial estimation of $g_{0}(\mathbf{a},W)$ and $\overline{Q}_{0}(\mathbf{a},W)$, 2. Evaluate $H_{\delta}(\mathbf{a}_{i},w_{i})$ for each observation, 3. Using the derived estimates, we implement a logistic regression to refine predictions under the shift $\text{logit}\overline{Q}_{\epsilon,n}(\mathbf{a}-\delta,w)=\text{logit}\overline{Q}_{n}(\mathbf{a}-\delta,w)+\epsilon H_{\delta,n}(\mathbf{a},w)$ then 4. From this model, once $\epsilon$ is determined, we update the predicted counterfactuals, using these updated counterfactuals, we then compute the plugin. This is a standard TMLE procedure for stochastic interventions; we simply shift multiple exposures simultaneously in the initial estimation and update. We estimate a standard error by taking the empirical variance of the estimated EIF and use this construct Wald-type confidence intervals and p-values. We use TMLE to ascertain unbiased estimates for each part of our interaction target parameter. That is, in $E_{\boldsymbol{\delta}}[Y]$ we update our counterfactual outcome under a joint shift by the joint density ratio as $H_{\delta}(\boldsymbol{a},w)$. Similarly, $E_{\delta_{1}}[Y]$ and $E_{\delta_{2}}[Y]$ are simply updated based on $H_{\delta}(a,w)$ for univariate shifts. In general, we use TMLE to target each of the three parts of our parameter. Given that our interaction is a continuous function (linear combination) of these asymptotically efficient estimates, we use the delta method for variance derivation.
@@ 4.1 Identifying Interactions in the Mixture Using Training Data
The discovery of important subsets of variables and interactions in the mixture variables is done using an ensemble method: the discrete Super Learner (SL) van der Laan Mark et al. [2007] and g-computation Robins [1987], Daniel et al. [2013]. The SL algorithm employs a cross-validation mechanism, optimizing over a library of candidate algorithms to select the estimator with the best empirical fit. We use a discrete SL to select the best fitting estimator from a library of different estimators for the estimation of $\mathbb{E}[Y|\boldsymbol{A},W]$. Using this SL, we then predict $\mathbb{E}[\mathbb{E}[Y|A_{i}-\delta_{i},\textbf{A}_{-i},W]]$, which is the individual exposure effects for exposure $A_{i}$, we do this for each exposure in the mixture. For all 2-way combinations, we then calculate: $\mathbb{E}[\mathbb{E}[Y|A_{i}-\delta_{i},A_{j}-\delta_{j},\mathbf{A}_{-i,j},W]]$, the expected outcome of a joint shift for all 2-way combinations. By comparing the estimated outcome under joint shift compared to the expected outcome sum of individual shifts for the same exposures and $\delta$’s we get a g-computation estimate for our interaction parameter, $\Psi_{\text{IE}}$, using the training data.
4.2 Ranking Synergistic and Antagonistic Interactions
For our purposes, we want to rank interaction effects based on their magnitude to identify the most significant synergistic and antagonistic effects, these define our oracle parameter sets of interactions:
$\mathcal{S}_{s}$ =$=\{(i,j)|\text{ranked by }\Psi_{\text{IE},ij}\text{ in descendingorder},\forall i<j\}[:s]$
$\mathcal{A}_{s}$ =$=\{(i,j)|\text{ranked by }\Psi_{\text{IE},ij}\text{ in ascending order},\forall i<j\}[:s]$
Where $\mathcal{S}_{s}$: are the top $s$ synergistic effects and $\mathcal{A}_{s}$ are the top $s$ antagonistic effects. This is because, if the joint impact is much greater than the additive impact for a change in exposures $i,j$, then this indicates a synergistic or superadditive relationship (the change in both is much more than additive). Inversely, if the interaction parameter is negative, this indicates that a change in both is less than the additive effect, which implies antagonism between the exposures.
In this framework, we have a family of parameters indexed by $(i,j)$, and we are particularly interested in the parameter $\Psi_{i_{0},j_{0}}(P)$ where $(i_{0},j_{0})$ maximizes the effect size over all possible pairs $(i,j)$. This maximizing parameter embodies the oracle parameter and helps us understand which interactions demonstrate the greatest synergy (most positive effect) or antagonism (most negative effect).
Defining these oracle parameters as the top synergistic/antagonistic interactions is imperative as they describe a consistent parameter that can be estimated in every fold of the CV procedure. That is, if we are simply interested in the max and min interaction effect (top synergistic and antagonistic interactions) this is estimable in every partition of the data, we can then pool our results for this parameter, and if the exposure sets used in the top ranks are consistent, we gain power for the interaction parameter and thus have tighter confidence intervals. Thus, our oracle target parameters are the top subsets of synergistic and antagonistic interactions based on our interaction parameter definition. The number of subsets, such as top 1 or top 3 ranked subsets are predefined by the user in our package.
@@ 4.3 Estimation of Shift Interventions on Discovered Variable Sets
Once these exposure sets for the top synergistic and antagonistic interactions are selected in the training data, we treat these exposure sets as fixed, we then estimate the resulting data-adaptively defined parameters using TMLE for marginal and joint stochastic shift interventions used to construct the interaction parameter. Using the training data, the nuisance functions $g_{n}$ and $Q_{n}$ are trained. The TMLE update step is performed using these training estimates to estimate $\epsilon$. With the validation data, we obtain estimates for $g_{n}$ and $Q_{n}$ and apply $\epsilon$ to debias our initial plug-in for our estimated stochastic shift target parameter using the validation data. In this case, the validation data are data that correspond to $k=1$. This procedure describes CV-TMLE Hubbard et al. [2018, 2016] which is the infrastructure we use to estimate our target parameter. Readers interested in the details of this procedure are directed to these resources.
@@
With the estimates for our parameter complete for the validation for $k=1$, we would then move on to $k=2$ and repeat the sections 4.1-4.3. Upon completion of the cross-estimation procedure, a pooled TMLE update provides a summary measure of the target parameter across k-folds. Specifically, we stack the predictions for each nuisance parameter from the estimation samples and run a pooled TMLE update on the cumulative initial estimates within each interaction rank (for example, the top synergistic interaction across all the folds). The resulting average is then used for parameter estimation. The pooled estimate across the folds can be succinctly represented as:
$\theta=\frac{1}{K}\sum_{k}\Psi_{F_{T_{n,k}}}\left(V_{n,k}\right)(P_{0})$
Where $\Psi_{F_{T_{n,k}}}$ identifies optimal exposure sets and subsequently formulates a plug-in estimator using training data $T_{n,k}$. $V_{n,k}$ refers to the estimation sample of the data for deriving our estimates.
@@ 4.4 Pooled Estimates under Data-Adaptive Delta
In stochastic interventions, the selection of shift magnitude $\boldsymbol{\delta}$ is critical, especially in the presence of joint stochastic shifts. These shifts can lead to positivity violations, which may bias and inflate the variance of the estimated exposure effects. To mitigate these risks, we employ a data-adaptive strategy where $\delta$ is dynamically adjusted within each training fold to maintain the ratio $H_{\delta}(a,w)_{i}=g_{n_{-k}}(a_{n_{-k}}-\delta\mid w)/g_{n_{-k}}(a_{n_{-k}}\mid w)$ below a pre-specified threshold $\lambda$ (default is 50 in our implementation). This ensures that all shifts are supported by the data, minimizing the risk of extreme or unsupported extrapolations.
When $\delta$ varies across folds due to differences in the data distributions or density estimates, we face the challenge of pooling results to form a coherent overall estimate. Instead of selecting a single $\delta$ such as the minimum or maximum, we opt to average the deltas. This approach is taken to reflect a shift scenario that is, on average, supported across all subsets of the data, thereby representing a more robust and generalizable intervention scenario. We then stack the nuisance parameters corresponding to these varied $\delta$ values and perform a TMLE update across the entire pooled dataset. We then plug these debiased counterfactuals into our influence function using the full data across various shifts. The resulting estimator reflects the average causal effect under these averaged shift scenarios.
This method acknowledges that, while individual fold-specific deltas might capture local data characteristics, the averaged delta provides a balanced view that is likely to be more stable and representative of the population as a whole. Nevertheless, this approach introduces a layer of complexity in interpretation, as the estimated effects correspond to an ’average’ intervention rather than a specific fixed intervention.
5 Density Estimation of the Exposures
Estimating the density of each exposure is crucial for evaluating the effects of interventions both on average across the population and within subpopulations. We provide two approaches to density estimation: direct density estimation and reparameterization based on classification, depending on user input.
5.1 Direct Density Estimation
Super Learner Algorithm: The Super Learner algorithm operates by training a variety of base learners (individual algorithms) on the data and then combining their predictions using a meta-learner. The meta-learner, in this case, is a non-negative least squares (NNLS) algorithm, which assigns weights to each base learner based on their performance. This approach leverages the strengths of different algorithms to produce a more accurate and robust predictive model.
Homoscedastic Errors (HOSE) Learner: The HOSE learner assumes that the variance of the errors is constant (homoscedasticity) across the range of predicted values. It employs a semi-parametric approach to density estimation, where the mean of the distribution is estimated using a Super Learner composed of multiple base learners.
Heteroscedastic Errors (HESE) Learner: The HESE learner, in contrast, allows for varying error variances (heteroscedasticity) across different levels of predicted values. It employs a two-step semi-parametric approach. The mean of the distribution is estimated using a similar Super Learner as in HOSE. An additional Super Learner is used to model the variance of the errors, allowing the error variance to change depending on the value of the predictors.
Although direct estimation is straightforward in principle, it can be challenging in practice due to the scarcity of effective data-adaptive estimators for high-dimensional conditional densities in the machine learning and statistical literature. As such, we also provide an alternative approach.
5.2 Reparameterization of Density Estimation via Classification
As an alternative to direct density estimation, we propose a novel approach that redefines the problem of estimating the density ratio $H(\mathbf{A},W)$ as a classification problem in an augmented dataset. This method leverages the fact that classification algorithms are well-developed and widely available in the machine learning literature compared to density estimators. This approach is used in other methodologies such as in longitudinal modified treatment policies Díaz et al. [2023].
To apply this approach, we construct an augmented dataset consisting of duplicated observations. For each original observation, we create a duplicate with the observed exposures $\mathbf{A}$ and another with the exposures under the intervention $\mathbf{A}^{\delta}$. We introduce an indicator variable $\xi$, which is set to 1 if the duplicated observation corresponds to the exposure under intervention, and 0 otherwise. Here, exposure can be the intervention on multiple exposures, such as the joint shift we need in our interaction parameter. The augmented dataset can be represented as:
$\{(W_{\xi,i,t},\mathbf{A}_{\xi,i,t},\xi_{\xi,i}):\xi=0,1;i=1,\ldots,n\},$
where $\xi_{\xi,i}=\xi$ indexes the duplicates, $W_{\xi,i,t}=W_{i,t}$ is the history variable, and $\mathbf{A}_{\xi,i,t}=\xi\cdot\mathbf{A}^{\delta}_{i,t}+(1-\xi)\cdot\mathbf{A}_{i,t}$ is the natural exposure level if $\xi=0$, and the intervened exposure level if $\xi=1$.
We denote the probability distribution of $(W_{t},\mathbf{A}_{t},\xi)$ in the augmented dataset by $P_{\xi}$, with corresponding density $p_{\xi}$. We define the parameter $u_{\xi}$ of $P_{\xi}$ as:
$u_{\xi}(\mathbf{a}_{t},w_{t})=P_{\xi}(\xi=1\mid\mathbf{A}_{t}=\mathbf{a}_{t},W_{t}=w_{t}).$
The density ratio $H(\mathbf{A},W)$ can be expressed in terms of $u_{\xi}$ as follows:
$H_{t}(\mathbf{a}_{t},w_{t})=\frac{p_{\xi}(\mathbf{a}_{t},w_{t}\mid\xi=1)}{p_{\xi}(\mathbf{a}_{t},w_{t}\mid\xi=0)}=\frac{P_{\xi}(\xi=1\mid\mathbf{A}_{t}=\mathbf{a}_{t},W_{t}=w_{t})}{P_{\xi}(\xi=0\mid\mathbf{A}_{t}=\mathbf{a}_{t},W_{t}=w_{t})}=\frac{u_{\xi}(\mathbf{a}_{t},w_{t})}{1-u_{\xi}(\mathbf{a}_{t},w_{t})},$
where the first equality follows from the definition of $H_{t}$ and the definition of conditional density, the second from Bayes’ rule and the observation that $P_{\xi}(\xi=1)=P_{\xi}(\xi=0)=\frac{1}{2}$, and the third from the definition of $u_{\xi}$.
By estimating $u_{\xi}$ in the augmented dataset using any classification method available in the machine learning and statistical learning literature (e.g., Super Learner), we can directly estimate the density ratio. This classification-based approach ensures that the estimator retains desirable properties such as asymptotic normality, provided that appropriate cross-validation and cross-fitting techniques are employed. Furthermore, this reclassification approach is substantially faster computationally, which is necessary, especially in our proposed methodology, where the conditional density is needed for each exposure and higher folds in cross-validation provide more consistent results. Overall, this approach has been used in previous literature to avoid density estimation. We simply create a duplicate data set, shift exposures by $\delta$ and create an intervention outcome = 1 for this data set. Bind rows with our unperturbed data with the outcome intervention = 0. Then we train a classifier and use Bayes rule to estimate the density ratio comparing shift to no shift.
In our implementation, users can choose between direct density estimation and the reparameterization approach via a user-defined parameter. Moving forward, all applications and simulations use the classification approach.
@@ 6 Simulations
In this section, we present simulations to demonstrate the effectiveness of our proposed method in identifying the correct exposures and interactions used to generate the outcome. The data-generating process (DGP) is designed to reflect the complexity of mixed exposures, where mixture variables are correlated but only some have a significant impact on the outcome.
6.1 Data-Generating Process
We consider a random vector of covariates, $\mathbf{W}=(W_{1},W_{2},W_{3},W_{4},W_{5},W_{6},W_{7})$, generated from a multivariate normal distribution with mean vector $\boldsymbol{\mu}=(0,0,0,0,0,0,0)$ and a covariance matrix $\boldsymbol{\Sigma}_{W}$. The covariance matrix $\boldsymbol{\Sigma}_{W}$ is designed to reflect real-world dependencies among covariates such as age, sex, BMI, systolic blood pressure (SBP), diastolic blood pressure (DBP), cholesterol, and fasting blood sugar (FBS).
Based on these covariates, we generate the exposures $\mathbf{A}=(M_{1},M_{2},M_{3},M_{4},M_{5},M_{6})$. Specifically, $M_{1}$ is a lognormal variable influenced by age ($W_{1}$) and sex ($W_{2}$): $M_{1}=\exp(0.2W_{1}+0.1W_{2}+\epsilon)$, where $\epsilon\sim\mathcal{N}(0,0.5)$. $M_{2}$ is a gamma variable influenced by BMI ($W_{3}$) and another variable ($W_{4}$): $M_{2}\sim\text{Gamma}(2,\exp(0.3W_{3}+0.2W_{4}))$. $M_{3}$ is a beta variable influenced by DBP ($W_{5}$) and random noise: $M_{3}\sim\text{Beta}(1+0.1W_{5},2+0.2\epsilon)$, where $\epsilon\sim\mathcal{N}(0,1)$. $M_{4}$ is a mixture of normal and uniform distributions (bimodal): $M_{4}=\mathbb{I}(\text{Uniform}(0,1)>0.7)\cdot\mathcal{N}(0.5W_{4},0.3)+\mathbb{I}(\text{Uniform}(0,1)\leq 0.7)\cdot\text{Uniform}(0,1)$. $M_{5}$ is a truncated normal variable influenced by cholesterol ($W_{6}$) and FBS ($W_{7}$): $M_{5}\sim\text{TruncNormal}(\text{min}(W_{6},\text{quantile}(W_{6},0.05)),\text{max}(W_{6},\text{quantile}(W_{6},0.95)),0.1W_{6}+0.2W_{7},0.4)$. $M_{6}$ is a mixture of normal distributions with right skew (lognormal-like): $M_{6}=\mathbb{I}(\text{Uniform}(0,1)>0.8)\cdot\mathcal{N}(0.2W_{5}+0.1W_{6},0.2)+\mathbb{I}(\text{Uniform}(0,1)\leq 0.8)\cdot(\mathcal{N}(0.2W_{5}+0.1W_{6},0.1)+\mathcal{N}(0,1))$.
The outcome variable $Y$ is then generated by incorporating both synergistic and antagonistic interactions among exposures: $Y=3+1.5M_{1}+1.2M_{2}-0.8M_{3}+2(M_{1}\times M_{2})+M_{4}+M_{5}+\epsilon$, where $\epsilon\sim\mathcal{N}(0,1)$ is a random noise term. Here, $M_{1}\times M_{2}$ represents a synergistic interaction that becomes stronger as the concentrations increase. Our goal is to accurately detect and estimate this synergistic interaction parameter as well as each part of it (the marginal and joint shifts used to construct the interaction parameter).
6.2 Analysis of Shifts and Interactions
We generate a large dataset $P_{0}$ of 100,000 units. For each pair of exposures, we apply shifts of 1 unit to each exposure and calculate the effects of single and pairwise shifts on the outcome. We compute the true difference between the combined effect of shifting both exposures and the sum of their individual effects to quantify the interactions. Specifically, we are interested in the following true effects built into the DGP: the top synergistic effect, which is 2.012, representing the average change in outcome beyond the additive effect (joint shift = 11.35 in $M_{1}$ and $M_{2}$ compared to the sum of individual shifts in these variables, $M_{1}=5.81$, $M_{2}=3.53$); and the top positive effect for a shift in $M_{1}$, which is 5.81.
These represent the true values we want to compare our estimates to. We run our InterXshift algorithm, which attempts to find the top synergistic, joint and marginal results. The goal is to assess the performance of our method in accurately identifying and estimating the impacts of synergistic interactions in a realistic mixed exposure scenario.